OpenStudy (anonymous):

(x^3+1/x^2+3x+2)(x^2-4/x^2-x+11)

OpenStudy (anonymous):

{(x+1)(x^2-x+1)/(x+1)(x+2)}{(x+2)(x-2)/??

OpenStudy (mimi_x3):

\[\large \frac{x^{3}+1}{x^{2}+3x+2} \times\frac{x^{2}-4}{x^{2}-x+11} \]?

OpenStudy (anonymous):

yes but I need to factor it out therefore my earlier reply

OpenStudy (mimi_x3):

\[\large \frac{(x+1)(x^{2}-x+1)}{(x+1)(x+2)} \times\frac{(x-2)(x+2)}{(x^{2}-x+11)} \] like that?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

x+1 cancel out so do x+2